# 3D Open-Vocabulary Panoptic Segmentation with 2D-3D Vision-Language Distillation

### v1

@@ Abstract
3D panoptic segmentation is a challenging perception task, which aims to predict both semantic and instance annotations for 3D points in a scene. Although prior 3D panoptic segmentation approaches have achieved great performance on closed-set benchmarks, generalizing to novel categories remains an open problem. For unseen object categories, 2D open-vocabulary segmentation has achieved promising results that solely rely on frozen CLIP backbones and ensembling multiple classification outputs. However, we find that simply extending these 2D models to 3D does not achieve good performance due to poor per-mask classification quality on novel categories. In this paper, we propose the first method to tackle 3D open-vocabulary panoptic segmentation. Our model takes advantage of the fusion between learnable LiDAR features and dense frozen vision CLIP features, using a single classification head to make predictions for both base and novel classes. To further improve the classification performance on novel classes and leverage the CLIP model, we propose two novel loss functions: object-level distillation loss and voxel-level distillation loss. Our experiments on the nuScenes and SemanticKITTI datasets show that our method outperforms strong baselines by a large margin.
@@ 1 Introduction
3D panoptic segmentation is a crucial task in computer vision with many real-world applications, most notably in autonomous driving. It combines 3D semantic and instance segmentation to produce per-point predictions for two different types of objects: things (e.g., car) and stuff (e.g., road). To date, there has been significant progress in 3D panoptic segmentation  [54, 44, 26, 49, 39, 40]. Most recently, methods such as [44] produce panoptic segmentation predictions directly from point clouds by leveraging learned queries to represent objects and Transformer [42]-based architectures [4, 2] to perform the modeling. However, existing models only predict panoptic segmentation results for a closed-set of objects. They fail to create predictions for the majority of unseen object categories in the scene, hindering the application of these algorithms to real-world scenarios.
@@
Figure 1: An illustration of open-vocabulary 3D panoptic segmentation results from our model. Without training on the categories of bus, trash can or vegetation, our method can produce accurate panoptic segmentation results even when the points are sparse.
@@
Recent advancements in vision-language (V-L) models [38, 21] have brought researchers’ attention to open-world computer vision tasks [3]. In 2D vision, there have been many successful methods in open-vocabulary object detection [14, 11, 23] and segmentation [10, 47, 51]. These methods learn to make predictions in a shared image-text embedding space, where predictions for unseen categories are produced by comparing the cosine similarity of an object with the text embedding of the category name. However, these methods are only possible due to the vast amounts of paired image-text data available, making it difficult to train similar models for 3D data such as LiDAR.
@@
Instead, researchers have continued to leverage the effectiveness of these 2D vision-language models for 3D with the help of pixel-point correspondences by running inference on 2D images and then aligning with the 3D features. These methods have achieved promising results on open-vocabulary semantic segmentation [9, 34, 50, 52] and instance segmentation [9, 50], individually. However, there are no methods that address the problem of 3D open-vocabulary panoptic segmentation, i.e., addressing both open-vocabulary semantic segmentation and open-vocabulary instance segmentation at the same time. The challenge lies in how to handle segmentation for novel things and stuff simultaneously.
@@
3D open-vocabulary panoptic segmentation is a challenging problem, both due to the significant domain gaps between the camera and LiDAR modalities and existing unsolved problems in open vocabulary segmentation. Many existing open-vocabulary works rely on similarities between text embeddings of class names and pre-trained V-L features to obtain associations between predictions and classes. However, while projecting 2D V-L features to 3D can account for a large part of the scene, there are often many points unaccounted for due to unmatched pixel / point distributions and differing fields of view between sensors. Many prior works rely entirely on the V-L features to perform both mask generation and object classification. While they work reasonably well, V-L models such as CLIP are trained at the image level, and so often suffer when tasked with fine-grained point or pixel level tasks.
@@
In this work, we aim to address these two issues with a novel architecture for 3D Open-Vocabulary Panoptic Segmentation. Building on existing 3D closed-set panoptic segmentation methods, we train a learned LiDAR feature encoder in parallel with a frozen, pre-trained camera CLIP model. By fusing the 3D LiDAR features with the 2D CLIP features, our model is able to learn rich features throughout the entire 3D sensing volume, even if there are no camera features in certain regions. In addition, we apply a pair of novel distillation losses that allow the 3D encoder to learn both object-level and voxel-level features which live inside the CLIP feature space. This provides a learned module in 3D space which can directly be compared with text embeddings. In addition, these losses also provide useful training supervision to unknown parts of the scene where there would otherwise be no loss gradient.
@@
With the proposed model and loss functions, our method significantly outperforms the strong baseline on multiple datasets. Our contributions are summarized as follows:
@@
  • We present the first framework for open-vocabulary 3D panoptic segmentation.
@@
  • We propose two novel loss functions, object-level distillation loss and voxel-level distillation loss to help novel object detection for both things and stuff classes.
@@ 2 Related Work
This work is closely related to 3D panoptic segmentation, 2D open-vocabulary segmentation, and 3D open-vocabulary segmentation and detection.
@@
2D open-vocabulary segmentation. 2D open vocabulary segmentation aims to group image pixels according to their semantics or identities for base (seen) or novel (unseen) categories. The prediction on novel categories is usually done by leveraging large V-L models [38, 21]. There are many works that focus on open vocabulary semantic segmentation [25, 46, 30, 13, 48, 53, 33, 16, 55, 56, 28]. Some work has also explored open-vocabulary panoptic segmentation [10, 47, 37]. Recently, FC-CLIP [51] proposes a single-stage framework based on a frozen convolutional CLIP backbone [31, 38, 20] for 2D open-vocabulary panoptic segmentation that achieves state-of-the-art performance. However, due to the camera-LiDAR domain gap, our experiments show that simply extending this 2D method to 3D leads to poor performance.
@@
3D open-vocabulary segmentation. 3D open-vocabulary segmentation is less explored due to the lack of 3D point-to-text association. One common practice is to utilize V-L models and use 2D-3D pairings to obtain rich, structured information in 3D [15, 6, 9, 50, 52, 34, 17, 7]. Notably, CLIP2Scene[6] proposes a semantic-driven cross-modal contrastive learning framework. PLA [9] leverages images as a bridge and builds hierarchical 3D-caption pairs for contrastive learning. OpenScene [34] extracts per-pixel CLIP features using a pre-trained V-L model [25, 13] then derives dense 3D features by projecting 3D points onto image planes. One unpublished concurrent work, RegionPLC [50], utilize regional visual prompts to create dense captions and perform point-discriminative contrastive learning, which can be used for semantic segmentation or instance segmentation, individually. In contrast, our work does not rely on any captioning model or extra contrastive learning, only pre-trained CLIP features. Our model also simultaneously handles semantic and instance segmentation.
@@ 3.1 Problem Definition
In open-vocabulary panoptic segmentation, the models are trained on $C_{B}$ base categories. At test time, besides these $C_{B}$ base categories, the data will contain $C_{N}$ novel categories. Following the settings of prior work [14, 23, 51], we assume the availability of the name of the novel categories during inference, but the novel categories are not present in the training data and their names are not known. Note that we do not apply any prompt engineering, as this is not the focus of this paper. We follow OpenScene [34] and obtain the CLIP Text embedding for each category.
@@
As there is no existing work for the 3D open vocabulary panoptic segmentation task, a natural idea would be extending the 2D open vocabulary segmentation methods to build the 3D counterpart. We start with P3Former [44], a state-of-the-art transformer-based 3D closed-set panoptic segmentation model, and add the essential components to support open-vocabulary capability by following FC-CLIP [51], a 2D open-vocabulary segmentation model that achieves state-of-the-art performance on multiple datasets. However, we found that this simple extension leads to poor performance in our experiments, and in this work we propose several new features to improve the performance of our model. More implementation details for this baseline can be found in the supplementary material.
@@ 3.3 Loss Function
Closed-set panoptic segmentation models [44] are typically optimized with objective functions consisting of a classification loss $L_{cls}$ and a mask prediction loss $L_{mask}$. We follow P3Former [44] for these two losses: the classification loss $L_{cls}$ optimizes the focal loss [29] between the class predictions and the category labels, while the mask loss $L_{mask}$ optimizes the voxel-query classification loss. Besides the two standard loss functions, we propose two new losses to apply distillation from the CLIP model at different levels.
3.3.1 Object-Level Distillation Loss
Similar to previous methods, our method uses the cosine similarity between predicted class embeddings and class text CLIP embeddings to produce per-class similarity scores. However, the classification loss applied to  Eq. 2 only enforces similarity to known classes. In this work, we make the assumption that the frozen CLIP features are discriminative with respect to open-vocabulary classes and have good out-of-distribution generalization. With this in mind, we propose an additional training loss which forces our predicted object-level class embeddings to be similar to the CLIP embeddings within their corresponding masks after matching. Similar to [51], we utilize voxel vision CLIP features to get a meaningful embedding for each query $q$ by mask pooling the Vision CLIP features:
@@ 3.3.1 Object-Level Distillation Loss
3.3.2 Voxel-Level Distillation Loss
While the object-level distillation loss distills the per-object features from CLIP model, it does not provide any supervision for the mask prediction head, which would otherwise only receive supervision for known classes. We found this particularly problematic for unknown stuff classes, which tend to be more spread out and cover larger and more diverse parts of the scene. In addition, it is only being applied to queries with relatively accurate mask predictions in order to learn useful CLIP features. To target these issues, we propose the voxel-level distillation loss to explicitly learn voxel-level CLIP features, which does not depend on any labels and can be applied on all queries. In particular, the voxel-level distillation loss is defined as:
@@ 3.3.2 Voxel-Level Distillation Loss
Unlike the object-level distillation loss, which is only applied to queries with matched ground truth, this loss is applied to all predicted mask scores and queries. In our experiments, we found that this loss significantly improves performance on stuff categories in particular, likely as it does not require exact matches with the ground truth, which can be difficult for large stuff classes. However, this loss is still susceptible to noisy or low quality mask scores, and we found that larger weights for this loss can disrupt training.
Our final objective function can be written as:
@@ 4 Experiments
Following the state-of-the-art closed-set 3D panoptic segmentation work, we conduct experiments and ablation studies on the nuScenes [4] and SemanticKITTI [12, 2] datasets.
@@ 4.1 Experimental Setting
nuScenes. The nuScenes dataset [4] is a public benchmark for autonomous driving tasks. It consists of $1000$ run segments and is further divided into prescribed train/val/test splits. We use all key frames with panoptic labels in the training set ($28130$ frames) to train the model. Following the most recent state-of-the-art model P3Former [44], we evaluate the models on the validation set ($6019$ frames). There are a total of $16$ labeled semantic classes, including $10$ things class and $6$ stuff classes.
@@
Data split. Both the nuScenes and SemanticKITTI datasets do not provide official base and novel class splits. Following the state-of-the-art 3D open-vocabulary segmentation work [9, 50], we randomly split the classes into base and novel, while keeping the ratio between base and novel classes around $3:1$. For nuScenes, the number of class for base and novel split are $12$ and $4$ respectively, and this setting will be referred as B12/N4. For SemanticKITTI, the number of class for base and novel split are $14$ and $5$, and this setting will be referred as B14/N5. We use the same splits in the main comparison with prior methods, and provide the results of more variations in the ablation studies.

### v2

@@ Abstract
3D panoptic segmentation is a challenging perception task, especially in autonomous driving. It aims to predict both semantic and instance annotations for 3D points in a scene. Although prior 3D panoptic segmentation approaches have achieved great performance on closed-set benchmarks, generalizing these approaches to unseen things and unseen stuff categories remains an open problem. For unseen object categories, 2D open-vocabulary segmentation has achieved promising results that solely rely on frozen CLIP backbones and ensembling multiple classification outputs. However, we find that simply extending these 2D models to 3D does not guarantee good performance due to poor per-mask classification quality, especially for novel stuff categories. In this paper, we propose the first method to tackle 3D open-vocabulary panoptic segmentation. Our model takes advantage of the fusion between learnable LiDAR features and dense frozen vision CLIP features, using a single classification head to make predictions for both base and novel classes. To further improve the classification performance on novel classes and leverage the CLIP model, we propose two novel loss functions: object-level distillation loss and voxel-level distillation loss. Our experiments on the nuScenes and SemanticKITTI datasets show that our method outperforms the strong baseline by a large margin.
Keywords:
Autonomous driving 3D panoptic segmentation Vision-language
footnotetext: * Work done while at Waymo
@@ 1 Introduction
3D panoptic segmentation is a crucial task in computer vision with many real-world applications, most notably in autonomous driving. It combines 3D semantic and instance segmentation to produce per-point predictions for two different types of objects: things (e.g., car) and stuff (e.g., road). To date, there has been significant progress in 3D panoptic segmentation  [58, 42, 40, 27, 52, 47]. Most recently, methods such as [47] produce panoptic segmentation predictions directly from point clouds by leveraging learned queries to represent objects and Transformer-based [45] architectures [4, 2] to perform the modeling.
@@
Figure 1: An illustration of 3D open-vocabulary panoptic segmentation results from our model. Without training on the categories of bus, trash can or vegetation, our method can produce accurate panoptic segmentation results even when the points are sparse.
However, existing models only predict panoptic segmentation results for a closed-set of objects. They fail to create predictions for the majority of unseen object categories in the scene, hindering the application of these algorithms to real-world scenarios, especially for autonomous driving. In this work, we focus on segmenting unseen $things$ and unseen stuff objects in autonomous driving scenarios. Concretely, we follow [10, 53] and develop models under the open-vocabulary setting: we divide the object categories into base (seen) categories and novel (unseen) categories, and evaluate models that are only trained on base categories.
@@
Such open-world computer vision tasks [3] benefit from the recent advancements in vision-language (V-L) models [39, 22]. In 2D vision, there are many successful methods in open-vocabulary object detection [15, 12, 24] and segmentation [11, 50, 54]. These methods make predictions in a shared image-text embedding space, where predictions for unseen categories are produced by comparing the cosine similarity of an object with the text embedding of the category name. However, these methods are only possible due to the vast amounts of paired image-text data available, making it difficult to train similar models for 3D data such as LiDAR.
@@
Instead, researchers have continued to leverage the effectiveness of these 2D vision-language models for 3D with the help of pixel-point correspondences by running inference on 2D images and then aligning with the 3D features. These methods have achieved promising results on open-vocabulary semantic segmentation [10, 35, 53, 55] and instance segmentation [10, 43, 53], individually. However, there are no methods that address the problem of 3D open-vocabulary panoptic segmentation, i.e., addressing both open-vocabulary semantic segmentation and open-vocabulary instance segmentation at the same time. The challenge lies in how to handle segmentation for novel things and stuff objects simultaneously.
@@
3D open-vocabulary panoptic segmentation is a challenging problem, due to both the significant domain gaps between the camera and LiDAR modalities and existing unsolved problems in open-vocabulary segmentation. Many existing open-vocabulary works rely on similarities between text embeddings of class names and pre-trained V-L features to obtain associations between predictions and classes [35, 55, 43]. However, while projecting 2D V-L features to 3D can account for a large part of the scene, there are often many points unaccounted for due to unmatched pixel/point distributions and differing fields of view between sensors. Some 3D open-vocabulary works [10, 53] apply contrastive learning to obtain better association between language and points, but they require extra captioning models and do not address the difficulties of detecting novel stuff classes.
@@
In this work, we aim to address these two issues with a novel architecture for 3D open-vocabulary panoptic segmentation. Building on existing 3D closed-set panoptic segmentation methods, we train a learned LiDAR feature encoder in parallel with a frozen, pre-trained camera CLIP model. By fusing the 3D LiDAR features with the 2D CLIP features, our model is able to learn rich features throughout the entire 3D sensing volume, even if there are no camera features in certain regions. In addition, we apply a pair of novel distillation losses that allow the 3D encoder to learn both object-level and voxel-level features which live inside the CLIP feature space. This provides a learned module in 3D space which can directly be compared with text embeddings. These losses also provide useful training supervision to unknown parts of the scene where there would otherwise be no loss gradient.
@@
With the proposed model and loss functions, our method significantly outperforms the strong baseline on multiple datasets.
Our contributions are summarized as follows:
@@
  • We present the first approach for 3D open-vocabulary panoptic segmentation in autonomous driving.
@@
  • We propose two novel loss functions, object-level distillation loss and voxel-level distillation loss to help segment novel things and novel stuff objects.
@@ 2 Related Work
This work is closely related to 3D panoptic segmentation, 2D open-vocabulary segmentation, and 3D open-vocabulary segmentation.
@@
2D open-vocabulary segmentation. 2D open-vocabulary segmentation aims to group image pixels according to their semantics or identities for base (seen) or novel (unseen) categories. The prediction on novel categories is usually done by leveraging large V-L models [39, 22]. There are many works that focus on open vocabulary semantic segmentation [26, 49, 31, 14, 51, 56, 34, 17, 57, 59, 29]. Some work has also explored open-vocabulary panoptic segmentation [11, 50, 38]. Recently, FC-CLIP [54] proposes a single-stage framework based on a frozen convolutional CLIP backbone [32, 39, 21] for 2D open-vocabulary panoptic segmentation that achieves state-of-the-art performance. However, due to the camera-LiDAR domain gap, we show that simply extending it to 3D leads to poor performance.
3D open-vocabulary segmentation. 3D open-vocabulary segmentation is less explored due to the lack of 3D point-to-text association. One common practice is to utilize V-L models and use 2D-3D pairings to obtain rich, structured information in 3D [16, 41, 7, 10, 53, 55, 35, 18, 8, 43]. Notably, CLIP2Scene[7] proposes a semantic-driven cross-modal contrastive learning framework. PLA [10] leverages images as a bridge and builds hierarchical 3D-caption pairs for contrastive learning. OpenScene [35] extracts per-pixel CLIP features using a pre-trained V-L model [26, 14] then derives dense 3D features by projecting 3D points onto image planes. One concurrent work, RegionPLC [53], utilize regional visual prompts to create dense captions and perform point-discriminative contrastive learning, which is used for semantic segmentation or instance segmentation, individually. In contrast, our work does not rely on any captioning model or extra contrastive learning, but only depends on pre-trained CLIP features. Our model also simultaneously handles semantic and instance segmentation.
@@ 3.1 Problem Definition
In open-vocabulary panoptic segmentation, the models are trained on $C_{B}$ base(seen) categories. At test time, besides these $C_{B}$ base categories, the data will contain $C_{N}$ novel(unseen) categories. Following the settings of prior work [15, 24, 54], we assume the availability of the name of the novel categories during inference, but the novel categories are not present in the training data and their names are not known. Note that we do not apply any prompt engineering, as this is not the focus of this paper. We follow OpenScene [35] to obtain the CLIP text embedding for each category.
@@
Most of the previous 3D open-vocabulary works only address semantic segmentation [16, 41, 7, 10, 53, 55, 35, 18, 8] or instance segmentation [53, 43] separately, and there is no existing work for the 3D open-vocabulary panoptic segmentation task, which handles novel things and novel stuff objects simultaneously. A natural idea would be extending the 2D open vocabulary segmentation methods to build the 3D counterpart. We start with P3Former [47], a state-of-the-art transformer-based 3D closed-set panoptic segmentation model, and add the essential components to support open-vocabulary capability by following FC-CLIP [54], a 2D open-vocabulary segmentation model that achieves state-of-the-art performance on multiple datasets. However, we found that this simple extension leads to poor performance in our experiments, and in this work we propose several new features to improve the performance of our model. More implementation details for this baseline can be found in the supplementary material.
@@ 3.3 Loss Function
Closed-set panoptic segmentation models [47] are typically optimized with objective functions consisting of a classification loss $L_{cls}$ and a mask prediction loss $L_{mask}$. We follow P3Former [47] for these two losses: the classification loss $L_{cls}$ optimizes the focal loss [30] between the class predictions and the category labels, while the mask loss $L_{mask}$ optimizes the voxel-query classification loss. Besides the two standard loss functions, we propose two simple yet effective losses to apply distillation from the CLIP model at different levels.
Object-level distillation loss. Similar to previous methods [50, 54], our method uses the cosine similarity between predicted class embeddings and class text CLIP embeddings to produce classification scores. However, the classification loss applied to  Eq. 2 only enforces similarity to known classes. In this work, we make the assumption that the frozen CLIP features are discriminative with respect to open-vocabulary classes and have good out-of-distribution generalization. We propose an additional training loss which forces our predicted object-level class embeddings to be similar to the CLIP embeddings within their corresponding masks after matching. Similar to [54], we utilize voxel vision CLIP features to get an embedding for each query $q$ by mask pooling Vision CLIP features:
@@
Voxel-level distillation loss. While the object-level distillation loss distills the per-object features from CLIP model, it does not provide any supervision for the mask prediction head, which would otherwise only receive supervision for known classes. We found this particularly problematic for unknown stuff classes, which tend to be more spread out and cover larger and more diverse parts of the scene. In addition, it is only being applied to queries with relatively accurate mask predictions in order to learn useful CLIP features. To target these issues, we propose the voxel-level distillation loss to explicitly learn voxel-level CLIP features, which do not depend on any labels and can be applied on all queries. In particular, the voxel-level distillation loss is defined as:
@@
Unlike the object-level distillation loss, which is only applied to queries with matched ground truth, this loss is applied to all predicted mask scores and queries. In our experiments, we found that this loss significantly improves performance on novel stuff categories in particular, likely as it does not require exact matches with the ground truth, which can be difficult for large stuff classes. However, this loss is still susceptible to noisy or low quality mask scores, and we found that larger weights for this loss can disrupt training.
To summarize, $L_{O}$ helps get rid of the ensemble of classifiers in [15, 14, 24, 50, 54] and enables open-vocabulary ability with one trainable classifier. $L_{V}$ uses a scene-level representation represented by the embedding of all queries, while previous methods only consider object-level representation. Combining $L_{O}$ with $L_{V}$ enables segmenting novel things and novel stuff objects simultaneously, which makes us the first work to address the 3D open-vocabulary panoptic segmentation problem in autonomous driving. Our final objective function can be written as:
@@ 4 Experiments
Following the state-of-the-art closed-set 3D panoptic segmentation work [58, 42, 40, 27, 52, 47], we conduct experiments and ablation studies on the nuScenes [4] and SemanticKITTI [13, 2] datasets.
@@ 4.1 Experimental Setting
nuScenes. The nuScenes dataset [4] is a public benchmark for autonomous driving. It consists of $1000$ run segments and is further divided into prescribed train/val/test splits. We use all key frames with panoptic labels in the training set($28130$ frames) to train the model. Following the most recent state-of-the-art model P3Former [47], we evaluate the models on the validation set($6019$ frames). There are $16$ semantic classes, including $10$ things classes and $6$ stuff classes.
@@
Data split. Both the nuScenes and SemanticKITTI datasets do not provide official base and novel class splits. Following the state-of-the-art 3D open-vocabulary segmentation work [6, 10, 53], we randomly split the classes into base and novel, while keeping the ratio between base and novel classes around $3:1$. For nuScenes, the number of class for base and novel split are $12$ and $4$ respectively, and this setting will be referred as B12/N4. For SemanticKITTI, the number of class for base and novel split are $14$ and $5$, and this setting will be referred as B14/N5. We use the same splits in the main comparison with prior methods, and provide the results of more variations in the ablation studies and supplementary materials.
@@
4.2 P3Former-FC-CLIP Baseline
As a baseline for novel-class panoptic segmentation, we construct a model from a fusion of P3Former [47] and FC-CLIP [54]. This baseline will be called P3Former-FC-CLIP (PFC). The baseline model takes the frozen voxel vision CLIP features as input, and the final prediction is obtained by geometric ensembling[15, 14, 24, 50, 54] of the results from the classification head $f_{cls}$ and another frozen classifier based on the similarity between the average-pool class embedding $w_{q}$ and the CLIP text embedding. Following FC-CLIP [54], the same set of learnable queries were used to represent both things and stuff classes. In summary, this baseline provides a comparison against our proposed method without the multimodal feature fusion module, the unified segmentation head, and the distillation losses. More information on this baseline can be found in the supplementary material.
Figure 4: Open-vocabulary panoptic segmentation results from PFC and our method on nuScenes. PFC predicts inaccurate category and masks for the novel pedestrian (red), bus (yellow) and vegetation (green), while ours makes correct predictions.
4.3 Main Results
Since there are no existing methods for the 3D open-vocabulary panoptic segmentation task, we mainly compare with three methods to demonstrate the capability of our method: (1) the strong open-vocabulary baseline method PFC to fairly demonstrate the strength of our method, (2) the closed-set state-of-the-art 3D panoptic segmentation method P3Former to understand the headroom of our method, and (3) the open-set, zero-shot state-of-the-art method for 3D semantic segmentation, OpenScene [35]. Comparisons on the nuScenes and SemanticKITTI datasets are shown in  Tab. 1 and  Tab. 3.
Results on nuScenes dataset. Tab. 1 shows the performance comparison on the validation set of the nuScenes dataset. Our method significantly outperforms the strong baseline PFC by a large margin across all metrics. PFC works relatively well for the novel things classes, but performance on the novel stuff class collapses. This is likely because stuff classes tend to cover large parts of the scene, leading to diverse per-voxel CLIP features which may not be good representatives for their respective classes. For the novel stuff classes, PFC only achieves $0.5$ $PQ_{N}^{St}$, while our method achieves $35.2$. PFC can produce reasonable masks for the novel stuff classes (with $SQ_{N}^{St}$=$60.4$), but fails to produce high quality semantic predictions for the novel stuff class (with $RQ_{N}^{St}$= $0.8$). Fig. 4 shows a qualitative comparison between PFC and our method. With our proposed design, and especially the new loss functions, our method is able to achieve strong performance for both novel things and stuff classes.